# PIGEON: Predicting Image Geolocations

### 17 Dec 2023 v4

@@ Abstract
Planet-scale image geolocalization remains a challenging problem due to the diversity of images originating from anywhere in the world. Although approaches based on vision transformers have made significant progress in geolocalization accuracy, success in prior literature is constrained to narrow distributions of images of landmarks, and performance has not generalized to unseen places. We present a new geolocalization system that combines semantic geocell creation, multi-task contrastive pretraining, and a novel loss function. Additionally, our work is the first to perform retrieval over location clusters for guess refinements. We train two models for evaluations on street-level data and general-purpose image geolocalization; the first model, PIGEON, is trained on data from the game of Geoguessr and is capable of placing over 40% of its guesses within 25 kilometers of the target location globally. We also develop a bot and deploy PIGEON in a blind experiment against humans, ranking in the top 0.01% of players. We further challenge one of the world’s foremost professional Geoguessr players to a series of six matches with millions of viewers, winning all six games. Our second model, PIGEOTTO, differs in that it is trained on a dataset of images from Flickr and Wikipedia, achieving state-of-the-art results on a wide range of image geolocalization benchmarks, outperforming the previous SOTA by up to 7.7 percentage points on the city accuracy level and up to 38.8 percentage points on the country level. Our findings suggest that PIGEOTTO is the first image geolocalization model that effectively generalizes to unseen places and that our approach can pave the way for highly accurate, planet-scale image geolocalization systems. Our code is available on GitHub. [^1] The GitHub link has been redacted in this preprint.
@@ 3.1 Geocell creation
We first employ naive, rectangular geocells inspired by the S2 library which subdivides every geocell until roughly balanced class sizes are reached. In contrast to S2 partitioning, our rectangular geocells are not of equal geographic size, creating even more balanced classes.
@@ 3.2 Hierarchical image geolocalization using distance-based label smoothing
Our approach addresses this major limitation and improves upon prior work by sharing all parameters between multiple, implicit levels of geographic hierarchies. We achieve this through a new loss function that relates adjacent geocells to each other, biasing the label based on the haversine distance which calculates the distance between two points on the Earth’s surface. Given two points, $\textbf{p}_{1}=(\lambda_{1},\phi_{1})$ and $\textbf{p}_{2}=(\lambda_{2},\phi_{2})$ with longitude $\lambda$ and latitude $\phi$, we define the haversine distance $\text{Hav}(\textbf{p}_{1},\textbf{p}_{2})$ in kilometers as follows:
@@ 3.3 Contrastive pretraining for geolocalization
In our experiments, we add a linear layer on top of CLIP’s vision encoder to predict geocells. For model versions with multiple image inputs (i.e. four-image panorama for PIGEON), we average the embeddings of all images. Averaging embeddings resulted in superior performance compared to combining multiple embeddings via multi-head attention or additional transformer layers.
In Haas et al. (2023), the authors demonstrate that continuing the pretraining of CLIP using domain-specific, synthetic captions derived from caption templates improves the generalized zero-shot performance on image geolocalization tasks. We further improve upon their work through the continued pretraining of CLIP in a multi-task fashion.
To this end, we augment our training datasets with geographic, climate, and directional auxiliary data. This data is used to create synthetic captions for each image by sampling caption components from different category templates and concatenating them. For PIGEOTTO, we use caption components based on the location, climate, and traffic direction. Meanwhile, for PIGEON, the Street View metadata allows us to additionally infer the compass direction and the season. Examples of caption components inferred from image metadata include:
@@ 4.2 Street View evaluation with PIGEON
-- Freezing Last CLIP Layer 91.82 255.1 45.47 4,531
@@
-- Freezing Last CLIP Layer 4.84 39.86 78.98 94.76 98.48
@@
For our final evaluation, we challenge one of the world’s foremost professional Geoguessr players to a match and win six out of six planet-scale, multi-round games. [^3] Link redacted for anonymity. PIGEON is the first model to reliably beat a Geoguessr professional.
@@ B. Implementation details
C. Auxiliary data sources
D. Ablation studies on non-distance metrics
E. Additional analyses
F. Deployment to Geoguessr
@@ Appendix A Semantic geocell creation
Algorithm 1 shows a slightly simplified version of how we split large geocells into multiple smaller ones without the help of administrative boundary information, resulting in a much more balanced geocell classification dataset. As one can see, the algorithm only depends on the geocell boundaries or shape definitions $g$, the training dataset $x$, an OPTICS clustering algorithm with parameters $p$ (can have round-specific parameters $p_{j}$), and a minimum cell size MINSIZE. The VORONOI algorithm takes a set of points as input and outputs a new geocell shape defined by these points which can be removed from the original cell shape.
@@ B.1 Model input
The biggest difference between PIGEON and PIGEOTTO is that PIGEON takes a four-image Street View panorama as input, whereas PIGEOTTO takes a single image as input. Images are always cropped to a square aspect ratio before being fed into the models. Figure 6 shows a representative input for PIGEON, depicting a 360-degree, four-image Street View panorama from a location in Pegswood, England.
Figure 6: Four images comprising a 360-degree panorama in Pegswood, England in our dataset.
PIGEOTTO’s training dataset is vastly different to PIGEON’s Street View input; the model takes a single image as input and was trained on a highly diverse image geolocalization dataset. Figure 7 shows eight images sampled from the MediaEval 2016 dataset Larson et al. (2017) which was derived from user-uploaded Flickr images. It is clearly visible that some of the images are extremely difficult to geolocalize, for example because they were taken indoors.
@@ B.2 Pretraining
Table 4 shows the hyperparameter settings used for the contrastive pretraining of CLIP for image geolocalization tasks. The CLIP weights were initialized with the pretrained weights of OpenAI’s CLIP implementation. [^4] https://huggingface.co/openai/clip-vit-large-patch14-336.
@@ B.3 Fine-tuning
Batch Size 256 256
@@ B.4 Hierarchical refinement
Appendix C Auxiliary data sources
@@
We obtain data on global climate zones through the Köppen-Geiger climate classification system Beck et al. (2018), with the data available here. Our planet-scale climate zone data is visualized in Figure 8. We use climate zone data both for synthetic caption generation for pretraining but also employ it when experimenting with multi-task prediction heads. In the latter case, climate zone prediction becomes a classification task.
@@
We obtain data on elevation through the United States Geological Survey’s Earth Resources Observation and Science (EROS) Center. As elevation data was missing for several locations in our dataset, we further augmented our data with missing values from parts of Alaska and parts of Europe, with the data for Alaska available here and the data for Europe available here. We use elevation data exclusively in a multi-task prediction setting via a log-transformed regression.
@@
We obtain data on average temperature, annual temperature range, average precipitation, and annual precipitation range through WorldClim 2. Similarly, the data is used in a multi-task regression setup, however, temperature values are not log-transformed before training.
@@
We obtain our driving side of the road data through WorldStandards. This data is used exclusively to generate synthetic captions for model pretraining.
Appendix D Ablation studies on non-distance metrics
Beyond the distance-based analysis of PIGEON described in the body of the paper, we also run ablation studies on non-distance metrics related to auxiliary data described in Appendix C. In Table 7, we observe that our final PIGEON model version actually does not perform best on non-distance metrics related to a location’s elevation, population density, season, and climate. The reason for this is that PIGEON does not share trainable model weights between the multi-task prediction heads and the location prediction tasks because joint multi-task training was already performed implicitly at the pre-training stage via synthetic captions. When sharing parameters between prediction heads (ablating “Freezing Last Clip Layer"), a positive transfer between the tasks is observed and better performances are achieved on these auxiliary prediction tasks.
A key takeaway from Table 7 remains that geographical, climate, demographic, and geological features can all be inferred from Street View images with potential applications in climate research.
@@ Appendix D Ablation studies on non-distance metrics
-- Freezing Last CLIP Layer 132.8 1,072 1.18 12.82 50.64 75.76
@@
Appendix E Additional analyses
E.1 Urban vs. rural performance
@@ Appendix E Additional analyses
E.2 Attention attribution examples
(a) Attention attribution map for an image in Canada.
(b) Attention attribution map for an image in New Zealand.
Contrastive pretraining used by CLIP gives the model a deeper semantic understanding of scenes and thereby enables it to discover strategies that are interpretable by humans. As we realized, the model was able to learn strategies that are taught in online Geoguessr guides without ever having been directly supervised to learn these strategies.
E.3 Failure cases
In spite of our model’s generally high accuracy of estimating image geolocations, there were several scenarios in which our model underperformed. By computing the entropy over the probabilities of all geocells for each location in our validation set, we managed to identify the images about which our model was the most uncertain. For PIGEOTTO, these were almost exclusively corrupted images remaining in the original Flickr training corpus. For PIGEON, however, which was solely trained on Street View images, we can observe some interesting failure cases in Figure 11.
@@ E.3 Failure cases
The features of poorly classified images are aligned with our intuitions and prior literature about difficult settings for image geolocation. Figure 11 shows that images from tunnels, bodies of water, poorly illuminated areas, forests, indoor areas, and soccer stadiums are amongst the imagery that is the most difficult to pinpoint geographically for a model trained on Street View data.
Appendix F Deployment to Geoguessr

### 07 Apr 2024 v5

@@ Abstract
Planet-scale image geolocalization remains a challenging problem due to the diversity of images originating from anywhere in the world. Although approaches based on vision transformers have made significant progress in geolocalization accuracy, success in prior literature is constrained to narrow distributions of images of landmarks, and performance has not generalized to unseen places. We present a new geolocalization system that combines semantic geocell creation, multi-task contrastive pretraining, and a novel loss function. Additionally, our work is the first to perform retrieval over location clusters for guess refinements. We train two models for evaluations on street-level data and general-purpose image geolocalization; the first model, PIGEON, is trained on data from the game of GeoGuessr and is capable of placing over 40% of its guesses within 25 kilometers of the target location globally. We also develop a bot and deploy PIGEON in a blind experiment against humans, ranking in the top 0.01% of players. We further challenge one of the world’s foremost professional GeoGuessr players to a series of six matches with millions of viewers, winning all six games. Our second model, PIGEOTTO, differs in that it is trained on a dataset of images from Flickr and Wikipedia, achieving state-of-the-art results on a wide range of image geolocalization benchmarks, outperforming the previous SOTA by up to 7.7 percentage points on the city accuracy level and up to 38.8 percentage points on the country level. Our findings suggest that PIGEOTTO is the first image geolocalization model that effectively generalizes to unseen places and that our approach can pave the way for highly accurate, planet-scale image geolocalization systems. Our code is available on GitHub. [^1] https://github.com/LukasHaas/PIGEON.
@@ 3.1 Geocell creation
We first employ naive, rectangular geocells inspired by the S2 library, subdividing every geocell until roughly balanced class sizes are reached. In contrast to S2 partitioning, our rectangular geocells are not of equal geographic size, creating even more balanced classes.
@@ 3.2 Hierarchical image geolocalization using distance-based label smoothing
Our approach addresses this major limitation and improves upon prior work by sharing all parameters between multiple, implicit levels of geographic hierarchies. We achieve this through a new loss function that relates adjacent geocells to each other, biasing the label based on the haversine distance which calculates the distance between two points on the Earth’s surface in kilometers. Given two points, $\textbf{p}_{1}=(\lambda_{1},\phi_{1})$ and $\textbf{p}_{2}=(\lambda_{2},\phi_{2})$ with longitude $\lambda$ and latitude $\phi$, and the earth’s radius $r$ in kilometers, we define the haversine distance $\text{Hav}(\textbf{p}_{1},\textbf{p}_{2})$ as follows:
@@ 3.3 Contrastive pretraining for geolocalization
In our experiments, we add a linear layer on top of CLIP’s vision encoder to predict geocells. For model versions with multiple image inputs (i.e. four-image panorama for PIGEON), we average the embeddings of all images. Averaging embeddings resulted in a superior performance compared to combining multiple embeddings via multi-head attention or additional transformer layers.
In Haas et al. (2023), the authors demonstrate that continuing the pretraining of CLIP using domain-specific, synthetic captions derived from caption templates improves the generalized zero-shot performance on image geolocalization tasks. We further improve upon their method through the continued pretraining of CLIP in a multi-task fashion.
To this end, we augment our training datasets with geographic, climate, and directional auxiliary data, used to create synthetic image captions by sampling caption components from different category templates and concatenating them. For PIGEOTTO, we use caption components based on the location, climate, and traffic direction. Meanwhile, for PIGEON, the Street View metadata allows us to additionally infer compass directions and the season, the latter included to avoid shortcut learning geirhos_et_al_2020 (i.e. snow $\rightarrow$ polar latitudes). Examples of caption components include:
@@ 4.2 Street View evaluation with PIGEON
-- Freezing Last CLIP Layer After Pretraining 91.82 255.1 45.47 4,531
@@
-- Freezing Last CLIP Layer After Pretraining 4.84 39.86 78.98 94.76 98.48
@@
For our final evaluation, we challenge one of the world’s foremost professional GeoGuessr players to a match and win six out of six planet-scale, multi-round games. [^3] https://www.youtube.com/watch?v=ts5lPDV--cU. PIGEON is the first model to reliably beat a GeoGuessr professional.
@@ B. Implementation details
C. Ablation study on pretraining captions
D. Ablation study on training datasets
E. Auxiliary data sources
F. Ablation studies on non-distance metrics
G. Additional analyses
H. Deployment to Geoguessr
I. Acknowledgements
@@ Appendix A Semantic geocell creation
Algorithm 1 shows a slightly simplified version of how we split large geocells into multiple smaller ones without the help of administrative boundary information, resulting in a much more balanced geocell classification dataset. As one can see, the algorithm only depends on the geocell boundaries or shape definitions $g$, the training dataset $x$, an OPTICS clustering algorithm with parameters $p$ (optionally round-specific parameters $p_{j}$), and a minimum cell size MINSIZE. The VORONOI algorithm takes a set of points as input and outputs a new geocell shape defined by these points which can be removed from the original cell shape.
@@ B.1 Model input
The biggest difference between PIGEON and PIGEOTTO is that PIGEON takes a four-image Street View panorama as input, whereas PIGEOTTO takes a single image as input. Images are always cropped to a square aspect ratio before being fed into the models. Figure 6 shows a representative input for PIGEON, depicting a 360-degree, four-image Street View panorama taken in Pegswood, England.
Figure 6: Four images comprising a 360-degree panorama from a location in Pegswood, England, in our dataset.
PIGEOTTO’s training dataset is vastly different to PIGEON’s Street View input; the model takes a single image as input and was trained on a highly diverse image geolocalization dataset. Figure 7 shows eight images sampled from the MediaEval 2016 dataset Larson et al. (2017) which was derived from user-uploaded Flickr images. It is clearly apparent that some of the images are extremely difficult to geolocalize, for example because they were taken indoors.
@@ B.2 Pretraining
Table 4 shows the hyperparameter settings employed for our contrastive pretraining of CLIP for the task of image geolocalization. The CLIP weights were initialized with the pretrained weights of OpenAI’s CLIP implementation. [^4] https://huggingface.co/openai/clip-vit-large-patch14-336.
@@ B.3 Fine-tuning
Haversine Smoothing $\tau$ 75 65
Batch Size 1024 1024
@@ B.4 Hierarchical refinement
Appendix C Ablation study on pretraining captions
In Section 3.3, we describe a novel multi-task contrastive pretraining method for image geolocalization. The ablation in Table 7 shows that our pretraining reduces PIGEON’s median kilometer error significantly from 57.8 to 44.4 kilometers ($-23.3\%$) versus no pretraining as in Wu & Huang (2022). Our innovation is that we are the first to design a multi-modal and multi-task contrastive pretraining objective for CLIP through the use of synthetic captions, and further find that the multi-task component of our method is highly effective; we observe a positive transfer from the auxiliary tasks embedded in our captions to the task of geolocalization, reducing our median error from 49.4 to 44.4 kilometers ($-10.2\%$) compared to pretraining solely with location captions as in Haas et al. (2023). Our multi-task contrastive pretraining method is general enough that it could also be employed in other problem domains.
Table 7: Ablation study of CLIP pretraining captions for PIGEON on a holdout dataset of 5,000 Street View locations.
Ablation Median Error Distance (@ km)
km 1 km 25 km 200 km 750 km 2,500 km
PIGEON [location + auxiliary captions] 44.35 5.36 40.36 78.28 94.52 98.56
PIGEON [location captions as in Haas et al. (2023)] 49.37 4.62 38.46 77.10 94.34 98.48
PIGEON [no pretraining as in Wu & Huang (2022)] 57.80 4.48 36.18 74.88 93.24 98.04
Appendix D Ablation study on training datasets
Section 4.1 in the body of our paper describes the different datasets used to train PIGEON and PIGEOTTO. While PIGEON was purely trained on Street View imagery, the training dataset for PIGEOTTO contains a combination of 4,166,186 geo-tagged images from the MediaEval 2016 dataset Larson et al. (2017) and 340,579 images from the Google Landmarks v2 dataset Weyand et al. (2020). Prior works’ benchmark results, listed in Table 3, employ a diverse range of training datasets with the goal of building the best performing and robust image geolocalization models. Since the prior SOTA model Geodecoder Clark et al. (2023) was exclusively trained on the MediaEval 2016 dataset Larson et al. (2017), we include an additional training dataset ablation for PIGEOTTO in Table 8 to distinguish data selection from system design effects.
Table 8: Ablation study of PIGEOTTO’s Google Landmarks v2  Weyand et al. (2020) data (340k images) against prior SOTA on five benchmarks.
Benchmark Method Median Distance (@ km)
Error Street City Region Country Continent
km 1 km 25 km 200 km 750 km 2,500 km
IM2GPS Hays & Efros (2008) GeoDecoder Clark et al. (2023) $\sim$ 25 22.1 50.2 69.0 80.0 89.1
PIGEOTTO [ME16] 75.6 11.8 38.8 63.7 80.6 91.1
PIGEOTTO [ME16 + Landmarks] 70.5 14.8 40.9 63.3 82.3 91.1
IM2GPS3k Vo et al. (2017) GeoDecoder Clark et al. (2023) $>200$ 12.8 33.5 45.9 61.0 76.1
PIGEOTTO [ME16] 163.6 10.9 35.8 52.4 70.7 84.4
PIGEOTTO [ME16 + Landmarks] 147.3 11.3 36.7 53.8 72.4 85.3
YFCC4k Vo et al. (2017) GeoDecoder Clark et al. (2023) $\sim 750$ 10.3 24.4 33.9 50.0 68.7
PIGEOTTO [ME16] 418.8 9.5 22.5 38.8 60.7 76.9
PIGEOTTO [ME16 + Landmarks] 383.0 10.4 23.7 40.6 62.2 77.7
YFCC26k Müller-Budack et al. (2018) GeoDecoder Clark et al. (2023) $\sim 750$ 10.1 23.9 34.1 49.6 69.0
PIGEOTTO [ME16] 356.5 10.1 24.6 41.3 62.6 78.7
PIGEOTTO [ME16 + Landmarks] 333.3 10.5 25.8 42.7 63.2 79.0
GWS15k Clark et al. (2023) GeoDecoder Clark et al. (2023) $\sim$ 2,500 0.7 1.5 8.7 26.9 50.5
PIGEOTTO [ME16] 440.8 0.1 8.7 30.1 64.0 84.7
PIGEOTTO [ME16 + Landmarks] 415.4 0.7 9.2 31.2 65.7 85.1
In Table 8, we observe that even when trained using the same data (ME16 Larson et al. (2017)), PIGEOTTO outperforms the prior SOTA Geodecoder Clark et al. (2023) by a large margin on four out of five benchmarks. The improvements in benchmark results can largely be attributed to the end-to-end design of PIGEOTTO, not our final training data selection. Still, we find that including the 340,579 landmark images Weyand et al. (2020) improves our model’s performance across all benchmarks and distance metrics. We further note that both PIGEOTTO versions are also more robust than Clark et al. (2023)’s Geodecoder by almost an order of magnitude, reducing the median geolocalization error by more than $5x$ on the out-of-distribution (OOD) benchmark dataset GWS15k Clark et al. (2023). Given the benchmark and OOD results, PIGEOTTO is currently the only planet-scale image geolocalization model robust to location and image distribution shifts.
Appendix E Auxiliary data sources
@@
We obtain data on global climate zones through the Köppen-Geiger climate classification system Beck et al. (2018), visualized in Figure 8. We use climate zone data both for synthetic caption generation for pretraining but also employ it in PIGEON’s ablation study as a classification task (ablating "Multi-task Prediction Heads"), described in Tables 1 and 2. The final PIGEON and PIGEOTTO versions only use climate zone data as part of their CLIP pretraining captions.
@@
We obtain data on elevation through the United States Geological Survey’s Earth Resources Observation and Science (EROS) Center. As elevation data was missing for several locations in our dataset, we augmented our data with missing values from parts of Alaska [^5] http://stacks.stanford.edu/file/druid:sg962yb7367/data.zip and Europe [^6] https://land.copernicus.eu/imagery-in-situ/eu-dem/eu-dem-v1.1/view. We use elevation data exclusively in a multi-task prediction setting via a log-transformed regression.
@@
We obtain data on the average temperature, annual temperature range, average precipitation, and annual precipitation range through WorldClim 2. Similarly to prior auxiliary data, temperate and precipitation data is used in a multi-task regression setup, however, temperature values are not log-transformed before training.
@@
We obtain data on the traffic direction through WorldStandards. This data is exclusively employed in generating synthetic pretraining captions.
Appendix F Ablation studies on non-distance metrics
Beyond the distance-based analysis of PIGEON described in the body of the paper, we also run ablation studies on non-distance metrics related to auxiliary data described in Appendix E. In Table 9, we observe that our final PIGEON model version actually does not perform best on non-distance metrics related to a location’s elevation, population density, season, and climate. The reason for this is that PIGEON does not share trainable model weights between the multi-task prediction heads and the location prediction tasks because joint multi-task training was already performed implicitly at the pretraining stage via synthetic captions. When sharing parameters between prediction heads (ablating “Freezing Last Clip Layer"), a positive transfer between the tasks is observed and better performances are achieved on these auxiliary prediction tasks.
A key takeaway from Table 9 remains that geographical, climate, demographic, and geological features can all be inferred from Street View images with potential applications in climate research and related fields.
@@ Appendix F Ablation studies on non-distance metrics
-- Freezing Last CLIP Layer After Pretraining 132.8 1,072 1.18 12.82 50.64 75.76
@@
Appendix G Additional analyses
G.1 Attention attribution examples
(a) Attention attribution map for an image in Canada.
(b) Attention attribution map for an image in New Zealand.
The contrastive pretraining used in CLIP gives the model a deeper semantic understanding of scenes and thereby enables it to discover strategies that are interpretable by humans. We observe that our model was able to learn strategies that are taught in online GeoGuessr guides without ever having been directly supervised to learn these strategies.
G.2 Urban vs. rural performance
@@ G.1 Attention attribution examples
G.3 Qualitative analyses of failure cases
Despite our models’ generally high accuracy in estimating image geolocations, there were several scenarios where they failed. We assess situations where our models were most uncertain and also identify the types of images for which our models made incorrect predictions.
Uncertainty.
By computing the entropy over the probabilities of all geocells for each location in our validation set, we identified images where our models were most uncertain. For PIGEOTTO, these images were almost exclusively corrupted images remaining in the original Flickr corpus. For PIGEON, however, which was solely trained on Street View images, we observe some interesting failure cases in Figure 11. The features of poorly classified images are aligned with our intuitions and prior literature about difficult settings for image geolocation. Figure 11 shows that images from tunnels, bodies of water, poorly illuminated areas, forests, indoor areas, and soccer stadiums are amongst the cases that are the most difficult to pinpoint by PIGEON.
@@ G.3 Qualitative analyses of failure cases
Incorrect Predictions.
While for PIGEON, most failure cases are out-of-distribution (OOD) images that are atypical for Street View imagery (Figure 11), PIGEOTTO was trained to be highly robust to distribution shifts with the goal of being a general image geolocation system. To evaluate in what cases PIGEOTTO fails, we collected representative images from the YFCC26k Müller-Budack et al. (2018) test set and plotted PIGEOTTO’s predictions against the ground truth coordinates on a map in Figure 13. We observe that real-world images (in this case derived from Flickr) are highly diverse, containing both indoor and outdoor images, vast differences in image depths, blurry images, images of people, filters that have been applied, and photos taken at night.
PIGEOTTO performs astoundingly well across a wide range of conditions; it correctly identifies popular places within circa one kilometer, as demonstrated by the third image in Figure 13 of the Capilano Suspension Bridge near Vancouver and the fourth image taken around the Kathmandu Durbar Square in Nepal. Our model fails in situations where the image contains very little information about the location, such as the fifth image containing only a boye in the sea, the sixth image containing a water bottle, or images taken at night with almost no visible features such as the eighth and thirteenth image, albeit in the later, PIGEOTTO does correctly predict Europe from the fireworks alone. PIGEOTTO further seems to work surprisingly well in indoor scenarios, even when images are blurred, as is the first image. Other examples include the eleventh image where our model still predicts the country correctly and the last image showing a person drinking from a red cup, guessed correctly to within 13 kilometers of the correct location. Finally, while PIGEOTTO confounds the wine regions of Victoria, Australia and Marlborough, New Zealand, it still is able to make accurate predictions from the flora alone, as evidenced by the close-up image of leaves, correctly guessed to within 671 kilometers.
Appendix H Deployment to GeoGuessr